At the beginning of 2018, Better Corps accounting records had the following general ledger accounts and balances: BETTER CORRequired a. Record the transactions in the appropriate general ledger accounts. Record the amounts of revenue, expense, and dEvent Assets BETTER CORPORATION Accounting Equation for 2018 = Liabilities + Stockholders Equity Land Common Retained PayablAs of December 31, 2018, determine the total amount of assets, liabilities, and stockholders equity and present this informa
Answer #1

a)

BETTER CORPORATION
Accounting Equation for 2018
Events Assets = Liabilities + Retained Earnings Account Titles for Retained Earnings
Cash + Land = Notes Payable + Common Stock + Retained Earnings
Balance 1/1/2018 $23000 + 33000 = 13300 + 8300 + 34400
1 (11500) + 11500 = + +
2 38000 + = + 38000 +
3 77000 + = + + 77000 Revenue
4 (40700) + = + + (40700) Expenses
5 23000 + = 23000 + +
6 (11500) + = + + (11500) Dividend
7 + = + +
Totals $97300 + $44500 = $36300 + $46300 + $59200

Assets= Cash+ Land= $97300+44500= $141800

Stockholders' Equity= Common stock+ Retained earnings= $46300+59200= $105500

Assets = Liabilities + Stockholders' Equity
b. $141800 = $36300 + $105500
c. $141800 = $36300 + $105500

The amount of assets, liabilities and stockholders' equity will be equal on Dec 31, 2018 and Jan 1, 2019